OpenStudy (anonymous):

find the exact values of six trigonometric functions of theta if theta is in standard position and P is on the terminal side. P(4,-3)

OpenStudy (anonymous):



OpenStudy (anonymous):

you get the hypotenuse either by pythagoras \(h^2=3^2+4^2\implies h=5\) or by remembering a 3 - 4 - 5 right triangle

OpenStudy (anonymous):

then use the appropriate ratios for example \[\sin(\theta)=-\frac{3}{5}\] etc